The price of good A rises from $2 to $ 4, causing a decrease in demand for good B from 10 units to 5 units. Calculate the cross price elasticity of demand


1
Expert's answer
2021-05-31T11:07:38-0400

Let's calculate cross price elasticity of demand of good Bย when the price of good A changes:


"E_{BA}=\dfrac{\dfrac{\Delta Q_B}{Q_B}}{\dfrac{\Delta P_A}{P_A}}=\dfrac{\dfrac{5-10}{\dfrac{5+10}{2}}}{\dfrac{\$4-\$2}{\dfrac{\$4+\$2}{2}}}=-1.0"